OpenStudy (anonymous):

A business purchases a computer system for $3000. If the value of the system decreases at a rate of 15% per year, how much is the computer worth after 4 years?

OpenStudy (anonymous):

you multiply by .85 repeatedly, so 3000*.85^4 .85 because 100% = 1.00 and 15% decrease = 1.00 - .15 = .85

OpenStudy (anonymous):

Wait huh? :I

OpenStudy (anonymous):

what part is confusing, I'll try to make it make sense...

OpenStudy (anonymous):

3000 * .85 = 2550 (after one year) 2550 * .85 = 2167.5 (after two years) but look, 2550 = 3000*.85, so after the second year can be found by: (3000*.85)*.85 which equals 3000*.85^2 so after 4 years, the exponent becomes a 4

OpenStudy (anonymous):

Each year its value decrease by 15% of its original value. In 4 years, it will decrease by 4*15%=60%. That means it will cost just 40% of its original value. So, its value after 4 years is: (40/100)*3000=$1,200

OpenStudy (anonymous):

No, the answer is not 1200

OpenStudy (anonymous):

because it is not losing 15% of 3000 each year. the second year, you lose 15% of a smaller number.

OpenStudy (anonymous):

OMG, Stick with tonks :)

OpenStudy (anonymous):

By AnwarA's calculation, you would have no money after 8 years, but you can lose 15% for ever

OpenStudy (anonymous):

Aha, I'm still trying to figure the problem out myself

OpenStudy (anonymous):

Monie, does using .85 make sense?

OpenStudy (anonymous):

Yea it does

OpenStudy (anonymous):

okay, then do it out slowly instead of any shortcuts

OpenStudy (anonymous):

But why .85?

OpenStudy (anonymous):

oh, if there were no change, you would have 100% the next year. Since there is a 15% decrease, you have 15% less than 100%, or 85% next year. Then move the decimal over to turn a percent into something you can multiply with for .85

OpenStudy (anonymous):

I end up with 0.7225...Is that correct? .-.

OpenStudy (anonymous):

hmmm, 3000*.85 = year 1 (3000*.85)*.85 = year 2 ((3000*.85)*.85)*.85 = year 3 (((3000*.85)*.85)*.85)*.85 = year 4 = 3000*.85*.85*.85*.85 = ...

OpenStudy (anonymous):

do you see why each step works? What do you get?

OpenStudy (anonymous):

I get 1556.01875.... But I'm guessing it to be 1556? Correct?

OpenStudy (anonymous):

I think you should keep it to the penny, 1556.02

OpenStudy (anonymous):

yes, do you get the .85 and the repeated multiplying?

OpenStudy (anonymous):

Yes I do :). THank you so very much for helping me Tonks.

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

One more question if you don't mind? (:

OpenStudy (anonymous):

If it is a math problem, post it as a new thread